OpenStudy (waheguru):

Solve THis Using Exponent laws (8^2)^6 PLEASE HELP ME I REALLY NEEDE HELP FAST

OpenStudy (saifoo.khan):

\[\Huge 8^{2 \times 6}\]

OpenStudy (waheguru):

but doesnt the 8 also get timesde by the six

OpenStudy (jamesj):

\[ (8^2)^6 = 8^2 . 8^2 . 8^2 . 8^2 . 8^2 . 8^2 \ \text{ (six times) } \] \[ = 8^{2+2+2+2+2+2} \] \[ = 8^{2 \times 6} \] \[ = 8^{12} \]

OpenStudy (waheguru):

so 8x6 = 48^12

OpenStudy (jamesj):

what? Where did that come from?

OpenStudy (waheguru):

doesnt the 8 also get multiplyed by 8 or does it stay the same

OpenStudy (jamesj):

By definition \( 8^{12} \) means 8 multiplied by itself 12 times \[ 8^{12} = 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \] Does that answer your question?

OpenStudy (waheguru):

ohhhhhhhhhhhh

OpenStudy (jamesj):

In the same way, \( (8^2)^6 \) means \(8^2 \) multiplied by itself 6 times: \[ (8^2)^6 = (8^2) \times (8^2) \times (8^2) \times (8^2) \times (8^2) \times (8^2) \]

OpenStudy (jamesj):

That is why the general law is \[ (x^a)^b = x^{ab} \] For your problem then, you can see that we use that law to immediately conclude that \[ (8^2)^6 = 8^{12} \] But you also see from above why that law makes sense and why in this particular case, the result is indeed true.